OpenStudy (anonymous):

Exponential Growth Formula Word Problem: When the population growth of a certain city was studied, the population was 22,000. It was found that the population, P, grows with respect to time, t, in years by the exponential formula P = (22,000)(10^0.0163t). How long will it take for the city to double its population?

OpenStudy (alekos):

so you need to solve for t when P = 44,000 correct?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

My answer was 18.468 years. Is this correct?

OpenStudy (alekos):

so we have an equation \[22,000.10^{.0163t} = 44,000\]

OpenStudy (alekos):

happy with that?

OpenStudy (anonymous):

Yes, so is 18.468 correct?

OpenStudy (alekos):

how did you get that?

OpenStudy (anonymous):

44,000 = 22,000(10^0.0163t) Divide 22,000 both sides 2 = 10^0.0163t log2 = 0.0163t t=18.468 did i do this correctly?

OpenStudy (alekos):

yes, well done that's correct

OpenStudy (anonymous):

Thanks